Given the demand and supply equations: (2.5)

Qdx= 150- 20Px and Qsx= -30 + 50Px

Find

a) Equilibrium Price

b) B) if market price is Rs. 3.50, what will be the amount of surplus

units?


1
Expert's answer
2021-11-09T17:52:35-0500

"Qd=Qs"

"150-20P=-30+50P"

"P=2\frac{4}{7}" Equilibrium price "Q=98\frac{5}{7}"

b) "(3.5-2\frac{4}{7})*98\frac{5}{7}=95\frac{4.5}{7}"